OpenStudy (anonymous):

Simplify? 1. ( 3 / 5 y ^ 4 ) ^-2 2. ( 2 / 5 n ^9) ^2

OpenStudy (anonymous):

1. you take a power of -2 of the power of 4 in that case multiply the powers and get -8 (3/5)^-2 x y^-8

OpenStudy (anonymous):

\[(a \times b)^2 = a^2 \times b^2\]yes or no ?

OpenStudy (anonymous):

I understand the first part, but i dont get what you're asking me for the second.

OpenStudy (anonymous):

we actually have to terms in the parentheses, 3/5 and y^4 I believe you have to take the power of ^-2 separately on both of them

OpenStudy (anonymous):

What are the answers for both of them?

OpenStudy (anonymous):

the answer for the first one is this (3/5)^-2 x y^-8